OpenStudy (anonymous):

Which mixed number is equal to 6.75? A. B. C. D.

OpenStudy (anonymous):

a 6 75/1000

OpenStudy (anonymous):

b 6 5/7

OpenStudy (anonymous):

c 6 3/4

OpenStudy (anonymous):

d 67/5

OpenStudy (er.mohd.amir):

6.75 when remove zeros it become 675/100

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

Step 1 : Convert 6.75 into mixed number.\[6.75=6\frac{ 75 }{ 100 }\]Step 2 :Try to reduce the result. We noticed that both 75 and 100 are divisible by 25.\[6\frac{ 75 }{ 100 }=6\frac{ 3\times25 }{ 4\times25 }\]

OpenStudy (anonymous):

\[=6\frac{ 3\times\cancel{25} }{ 4\times\cancel{25} }\]\[=?\]solve it @mjmahmood

OpenStudy (anonymous):

75 and100

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@HelperJaje

OpenStudy (inarito):

Its c.

OpenStudy (inarito):